OpenStudy (anonymous):

what is the integration of tan^-1(1/(x^2-x+1))

OpenStudy (anonymous):

I think integration by parts is the way to go here. Let \[\begin{matrix} u=\tan^{-1}\left(\frac{1}{x^2-x+1}\right)&dv=dx\\ du=\cdots~~~~~~~~~~~~~~~~~~~~~~~~~&v=x \end{matrix}\] Completing the square in the denominator and making a substitution there might make finding \(du\) easier.

zepdrix (zepdrix):

Hmm yah I think sith has the right idea here. \[\large= x \arctan\left(\frac{1}{x^2-x+1}\right)-\int\limits x\frac{\left(\dfrac{1}{x^2-x+1}\right)'}{1+\left(\dfrac{1}{x^2-x+1}\right)^2}dx\] That new integral looks like a doozy to clean up.. but it's probably the right way to go.

OpenStudy (noelgreco):

Complete the square of the denominator and it should look familiar.

zepdrix (zepdrix):

\[\large x^2-x+1 \qquad = \qquad x^2-x+\left(-\frac{1}{2}\right)^2-\left(-\frac{1}{2}\right)^2+1\]\[\large =\left(x-\frac{1}{2}\right)^2+\frac{3}{4}\] @NoelGreco , it should look familiar? D: Hmm I'm not quite seeing it...

OpenStudy (anonymous):

\[\int\limits x\frac{\left(\dfrac{1}{x^2-x+1}\right)'}{1+\left(\dfrac{1}{x^2-x+1}\right)^2}dx=\int\limits -x\frac{\left(\dfrac{2x-1}{(x^2-x+1)^2}\right)}{1+\left(\dfrac{1}{x^2-x+1}\right)^2}dx\]\[\int \frac{-x(2x-1)}{1+(x^2-x+1)^2} dx\]how to do it from here?

OpenStudy (anonymous):

i can not understand the given process........sorry

OpenStudy (anonymous):

is there any one to help.............

OpenStudy (tkhunny):

Unless there is some really important reason to do otherwise, this looks like a candidate for numerical integration. I'd like to see the ENTIRE original problem statement.

OpenStudy (anonymous):

ok the original problem statement is to,,,intigrate....tan^-1(1/x^2-x+1)

OpenStudy (anonymous):

answer brought to you by Matlab. It doesn't show steps tho. arctan(x - 1) + ln(x^2 - 2*x + 2)/2 - ln(x^2 + 1)/2 + x*arctan(1/(x^2 - x + 1))

zepdrix (zepdrix):

\[\large \int\limits \frac{-x}{1+(x^2-x+1)^2} (2x-1)dx\] \(\large u=x^2-x+1\) \(\large du=(2x-1)dx\) \[\large \int\limits \frac{-x}{1+u^2}du\] Hmmmmm......

OpenStudy (anonymous):

wht is meant by arctan?@chemENGINEER

OpenStudy (anonymous):

how have u strted?????@zepdrix

OpenStudy (anonymous):

hey one thing @zepdrix how come we have x left in the integral if we are integrating w.r.t u?i mean it will then be treated as a constnt isn't it?

zepdrix (zepdrix):

No not constant, it needs to be dealt with. So there's an issue there :P that's why i stopped. Kinda stuck.

zepdrix (zepdrix):

@shaown I'm not sure if the start process was correct, so I'm not sure if it's worth explaining in detail.. hmm

zepdrix (zepdrix):

Are you familiar with `Integration by Parts` Shaown? :o

OpenStudy (experimentx):

\[ \arctan \left( \frac{1}{x^2 - x + 1} \right) = \arctan\left( \frac{x - (x-1)}{1 + x(x-1)}\right ) = \arctan x - \arctan(x-1) \] From here, use Integration by parts.

OpenStudy (anonymous):

neat as usual @experimentX